A transverse wave 0.80 cm in amplitude is propagating on a string; the wave frequency is 46 Hz. The string is under 21 N tension and has mass per unit length of 15 g/m. Determine the wave speed. Answer should be in m/s. Thanks!
Answer #1

\mu = mass per unit length = 15 g/m = 0.015 kg/m

wave speed = V = sqrt(T/\mu) = sqrt (21/0.015) = 37.42 m/s